OpenStudy (anonymous):

Using complete sentences, compare and contrast the "sum of two squares" and the "difference of two squares." Provide an explanation and example of each, including any similarities and differences.

OpenStudy (anonymous):

the sum of two squares provide a perfect trinomial, while difference of two squares provides perfect binomial. .

OpenStudy (anonymous):

sum of 2 squares can't be factored over the reals, difference of 2 squares can

OpenStudy (anonymous):

x^2-9=(x-3)(x+3) x^2+9=(x-3i)(x+3i)